If your uncle borrows $55,000 from the bank at 9 percent interest over the seven-year life of the loan. Use Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

 

What equal annual payments must be made to discharge the loan, plus pay the bank its required rate of interest? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
 




How much of his first payment will be applied to interest? To principal? (Do not round intermediate calculations. Round your final answers to 2 decimal places.)
 




 How much of his second payment will be applied to each? (Do not round intermediate calculations. Round your final answers to 2 decimal places.)
Answer #1

a. 

Equated annual payments 

= (P*r*(1+r)^n) / ((1+r)^n - 1)

= (55000*9%*1.09^7)/(1.09^7-1)

= 9048.79365/0.82804

= 10927.97


b. 

Opening balance = 55000


Interest aaccrued for the year = 55000*0.09 = 4950


Payment for the year = 10927.97


Hence interest component = 4950


Principle ccomponent = 10927.97-4950 = 5977.97


Closing balance = 55000 + 4950 - 10927.97 = 49052.03


c. 

Use the same method as above, interest component = 9%*49052.03 = 4414.6827


Hence principle component = 10927.97 - 4414.6827 = 6513.2873


answered by: gavin